OpenStudy (anonymous):

A good rule of thumb to follow when dealing with rational exponents that are negative is to make them positive before working through the problem. To change a term with a negative exponent to a term with a positive exponent you simply need to rewrite the term on the _____________ side of the fraction bar and change the sign to ______________.

zepdrix (zepdrix):

Hmm I'm not exactly sure what words they're looking for. I would guess: " To change a term with a negative exponent to a term with a positive exponent you simply need to rewrite the term on the `other` side of the fraction bar." Example:\[\large \frac{x^{-2}}{1} \qquad=\qquad \frac{1}{x^2}\] Or another example:\[\large \frac{1}{x^{-3}} \qquad=\qquad \frac{x^3}{1}\]

zepdrix (zepdrix):

"a term with a negative exponent to a term with a positive exponent ... and change the sign to ______________." So with the way they worded it, we're starting with a `negative exponent`. So if you had to change the sign of the exponent, what would it change to?

OpenStudy (anonymous):

positive